# On the Black-box Explainability of Object Detection Models for Safe and Trustworthy Industrial Applications

### v1

@@ Abstract
In the realm of human-machine interaction, artificial intelligence has become a powerful tool for accelerating data modeling tasks. Object detection methods have achieved outstanding results and are widely used in critical domains like autonomous driving and video surveillance. However, their adoption in high-risk applications, where errors may cause severe consequences, remains limited. Explainable Artificial Intelligence (XAI) methods aim to address this issue, but many existing techniques are model-specific and designed for classification tasks, making them less effective for object detection and difficult for non-specialists to interpret. In this work we focus on model-agnostic XAI methods for object detection models and propose D-MFPP, an extension of the Morphological Fragmental Perturbation Pyramid (MFPP), which uses segmentation-based mask generation. Additionally, we introduce D-Deletion, a novel metric combining faithfulness and localization, adapted specifically to meet the unique demands of object detectors. We evaluate these methods on real-world industrial and robotic datasets, examining the influence of parameters such as the number of masks, model size, and image resolution on the quality of explanations. Our experiments use single-stage object detection models applied to two safety-critical robotic environments: i) a shared human-robot workspace where safety is of paramount importance, and ii) an assembly area of battery kits, where safety is critical due to the potential for damage among high-risk components. Our findings evince that D-Deletion effectively gauges the performance of explanations when multiple elements of the same class appear in the same scene, while D-MFPP provides a promising alternative to D-RISE when fewer masks are used.
@@ 1 Introduction
In this context, the field of Explainable AI (XAI) Barredo Arrieta et al. (2020) aims to enhance the interpretability of AI systems by their audience and ultimately, to enhance the user’s trust in the output of AI-based systems. Leaving aside the category of transparent AI models (which are inherently interpretable and do not require any explanations for a user to understand how they work), explainability methods in XAI can be broadly categorized into white-box and black-box approaches. White-box XAI methods require access to the internal workings of the model, such as weights, activations, or gradients (e.g., Grad-CAM Selvaraju et al. (2017)). While these methods can provide powerful insights, they are often limited by their dependence on specific model architectures, making them difficult to generalize across different models and less accessible to users unfamiliar with AI research/tools. In contrast, black-box XAI methods treat the model as an opaque entity, providing explanations based solely on the model’s input-output behavior without requiring any access to its internal components. However, most black-box XAI methods are designed for classification tasks rather than for object detection Ribeiro et al. (2016); Lundberg and Lee (2017); Petsiuk (2018); Ali et al. (2023a).
@@
  • Last but not least, we facilitate the broader adoption of the developed techniques for object detection in real-world use cases by releasing the code publicly in a repository https://github.com/aklein1995?tab=repositories  [^1] To be published upon article acceptance..
@@ 2 Related Work
As stated in the introduction, XAI offers insights into the procedure followed by an AI-based system to elicit their outputs, enabling end-users to understand and eventually trust the decisions output by the AI-based system grounded on objective data Ali et al. (2023b). To date, the majority of XAI methods are designed for models learned to address classification tasks. For instance, CAM-based methods like GradCAM Selvaraju et al. (2017), GradCAM++Chattopadhay et al. (2018) and Integrated Gradients Sundararajan et al. (2017) quantify and attribute the pixel-wise importance of a given input according to the gradients with respect a target class. Moreover, making use of backpropagation, LRP Montavon et al. (2019) calculates the contribution that a neuron has with neurons in consecutive layers to get relevance scores. In contrast, perturbation-based techniques work by occluding certain parts of the input and analyzing its impact in the predictions. Within this type of techniques, LIME Ribeiro et al. (2016), approximates a NN with an interpretable model; SHAP Lundberg and Lee (2017) assigns importance values to each input feature based on Shapley values; RISE Petsiuk (2018) generates saliency maps by probing the model with randomly masked versions of the input image; and MFPP Yang et al. (2020) generates masks by dividing the input image into multi-scale superpixels. Nonetheless, none of them have been explicitly extended for object detection tasks –with the exception of RISE, which has been adapted for this purpose– although techniques like SHAP can also be utilized for regression problems.
@@ 3 Background
We now proceed by elaborating on key concepts needed to properly understand the details of the proposed D-MFPP technique and the D-Deletion metric that lie at the core of this work. Concretely, we provide fundamentals for object detection models (Subsection 3.1) and attribution-based for object detection
@@ 3.1 Object Detectors
Object detectors are crucial components in computer vision tasks, capable of identifying and localizing objects within an image. They can be broadly categorized into one-stage and two-stage detectors.
One-stage detectors
They directly predict bounding boxes and class probabilities from input images in a single pass. These detectors treat object detection as a simple regression problem, straight from image pixels to bounding box coordinates and class probabilities. To this end, they produce a dense grid of bounding box proposals and class probabilities in one step. Specifically, YOLO divides the input image into a grid and predicts bounding boxes and class probabilities for each grid cell. One-stage detectors are usually adopted for real-time applications due to their high speed and efficiency, although their accuracy tends to be lower when being compared with two-stage detectors.
@@
These models, among which Faster R-CNN Ren et al. (2017) can be considered to be the most representative one, follow a more complex approach that divides the detection process into two stages. In the first stage, a Region Proposal Network (RPN) generates a set of candidate object proposals (bounding boxes) from the input image. In the second stage, these proposals are refined and classified into different object categories by a second network. This second stage typically involves a more complex network, such as a convolutional neural network (CNN), which performs classification and further refinement of the bounding box coordinates. Two-stage detectors are known for their high accuracy and robustness in detecting objects of varying sizes and shapes, but they are generally slower than one-stage detectors.
@@
where $\mathbf{L}_{i}$ defines the bounding box corners $(x^{i}_{1},y^{i}_{1})$ and $(x^{i}_{2},y^{i}_{2})$; $O_{i}\in[0,1]$ refers to the probability that bounding box $L_{i}$ contains an object of any class; and $\mathbf{P}_{i}$ is a vector of probabilities $(p^{i}_{1},\ldots,p^{i}_{C})$ representing the probability that region $\mathbf{L}_{i}$ belongs to each of $C$ classes. As it can be observed in this above definition, unlike traditional classifiers which assign a single class label to an entire image, object detectors must simultaneously handle both classification and localization. This dual task involves predicting not only the class of each object but also its precise location within the image. The complexity of this task increases the difficulty of making these models interpretable.
@@ 3.2 Explainable Artificial Intelligence (XAI)
Despite the great performance exhibited by object detectors in manifold applications, their adoption in risk-sensitive scenarios is often hindered by a lack of trust and transparency by the user making decisions based on the detections issued by these models. As introduced previously, research on XAI produce techniques and methods that make the behavior and predictions of AI models understandable to humans without sacrificing performance Gunning et al. (2019). To this end, multiple XAI techniques have been proposed, which can be classified into four broad categories Ali et al. (2023b):
@@
Among black-box XAI methods, perturbation-based approaches are commonly used due to their simplicity and effectiveness in revealing which parts of the input are most influential for the model’s predictions. Perturbation-based techniques offer a direct way to assess how changes to the input image affect the model’s output. By systematically altering or masking parts of the input image (using masks to generate perturbed samples), these methods allow inferring the importance of different regions based on the model’s input-output behavior.
@@ 4 Materials and Methods
This section describes the industrial robotics use cases providing the datasets (Subsection 4.1), XAI methods (Subsection 4.3) and the explanation quality metrics (Subsection 4.4) considered in our work. The novel XAI technique and quality metrics proposed in this manuscript are also described in detail in Subsection 4.3.
@@ 4.1 Industrial Robotics Datasets under Consideration
The datasets used in this manuscript have been collected during the course of the ULTIMATE project, https://ultimate-project.eu/, which features two distinct real robotics use cases Kozik et al. (2024). The first dataset, from PIAP https://piap.lukasiewicz.gov.pl/), involves a collaborative workspace where a human and a robotic arm work together. The second dataset, provided by Robotnik https://robotnik.eu/, focuses on a battery assembly area, where a robotic arm assembles components for a battery kit.
@@ 4.3 Explainability Methods
We evaluate four popular methods for generating visual explanations of black-box models: LIME, RISE, D-RISE, and D-MFPP. The first two methods, LIME and RISE, were originally developed for image classifiers but can be adapted to object detectors. However, they primarily focus on explaining classification aspects and are not capable of addressing localization characteristics. In contrast, D-RISE is one of the first XAI methods specifically designed for object detectors, providing explanations that encompass both classification and localization. Additionally, we extend the existing MFPP method (originally tailored for classifiers) into a version suitable for object detection, which we refer to as D-MFPP. In what follows we briefly describe them, flowing into a description of the proposed D-MFPP approach:
@@ 4.4 Metrics
Evaluating the performance of attribution-based explainability methods for image data involves assessing how well the generated relevance heatmaps highlight important regions of the input image that contribute to the model’s decision. Generally, according to Hedström et al. (2023), explanation quality metrics can be grouped into six categories based on their logical similarity: faithfulness, robustness, localization, complexity, randomization, and axiomatic metrics. In this study, we focus on two of these categories that are particularly relevant to object detection: localization (Subsection 4.4.1) and faithfulness (Subsection 4.4.2).
@@ 4.4.2 Faithfulness
  • Minimum Subset: It follows the same logic as Deletion, but instead of determining the AUC, it considers the required number of pixels that make the prediction to change Gevaert et al. (2024). Given the importance score for each pixel ($S_{I,f}$), Min-Subset is defined as the smallest subset of pixels that needs to be removed to change the model’s prediction. Mathematically,
@@ 4.4.3 Proposed D-Deletion and D-Minimal Subset Metrics
The difference between Deletion and D-Deletion is illustrated in Figure 4. This figure highlights how D-Deletion distinguishes between different objects of the same class by incorporating localization information, leading to more refined and accurate explanations ($\downarrow$ AUC in the Figure’s last row) when multiple objects of the same class are detected in an image. For the sake of clarity, we provide the pseudocode of Deletion in Algorithm 1, where the main difference with respect to D-Deletion are lines 9 to 11.
@@ 5 Experiments and Results
  • RQ1. Which XAI method provides the most reliable and insightful explanations for object detection models?
  • RQ2. Does the D-Deletion metric enhance the trustworthiness of XAI outputs when multiple objects of the same class are present in the image?
  • RQ3. How does the mask generation process influence the quality of explanations, particularly when using similarity scores for object detection? How does D-MFPP behave?
  • RQ4. Do different image dimensions impact the explanations generated by XAI methods? Do models of varying sizes (large, medium, small, nano) focus on different regions of the image in their explanations?
@@
RQ1. Comparison between XAI methods
@@
Table 1: Quantitative metrics of LIME, RISE and D-RISE for the Human-Robot dataset. The table presents the performance of each XAI technique in terms of classification (Deletion, D-Deletion, Min-Subset, D-Min-Subset) and localization metrics (PG and EBPG), with scores reported for each class and the overall average. $\downarrow$ indicates that lower values are better, whereas $\uparrow$ indicate that higher values are better. The best results for each metric and object type are highlighted in gray, while the best average values are written in bold.
@@
Table 2: Quantitative metrics of LIME, RISE and D-RISE for the Battery Assembly dataset. The table presents the performance of each XAI technique in terms of classification (Deletion, D-Deletion, Min-Subset, D-Min-Subset) and localization metrics (PG and EBPG), with scores reported for each object and the overall average. The best results for each metric and object type are highlighted in gray, while the best average values are marked in bold.
@@
RQ2. D-Deletion metric for scenes with multiple objects of the same class
RQ3. Influence of the Mask Generation Strategy
Influence of the number of masks
Table 3: Quantitative results for saliency maps generated with various masks techniques (D-Sliding Window, D-RISE, and D-MFPP) in the Human-Robot dataset. Only the average score across all classes is presented. The best mask configuration for each XAI method is highlighted in gray, while the overall best-performing XAI method with its optimal mask configuration is shown in bold.
@@
Table 4: Quantitative results for saliency maps generated with various masks techniques (D-Sliding Window, D-RISE, and D-MFPP) in the Battery Assembly dataset. The best mask configuration for each XAI method is highlighted in gray, while the overall best-performing XAI method with its optimal mask configuration is shown in bold.
@@
RQ4. Image dimension variation and YOLOv8 complexity
Table 5: Explanations comparison for three image dimensions (720x1280, 360x640, and 180x320) and varying resolutions (16x16, 8x8, and 4x4) in the Human-Robot dataset.

### v2

@@ Abstract
In the realm of human-machine interaction, artificial intelligence has become a powerful tool for accelerating data modeling tasks. Object detection methods have achieved outstanding results and are widely used in critical domains like autonomous driving and video surveillance. However, their adoption in high-risk applications, where errors may cause severe consequences, remains limited. Explainable Artificial Intelligence methods aim to address this issue, but many existing techniques are model-specific and designed for classification tasks, making them less effective for object detection and difficult for non-specialists to interpret. In this work we focus on model-agnostic explainability methods for object detection models and propose D-MFPP, an extension of the Morphological Fragmental Perturbation Pyramid (MFPP) technique based on segmentation-based masks to generate explanations. Additionally, we introduce D-Deletion, a novel metric combining faithfulness and localization, adapted specifically to meet the unique demands of object detectors. We evaluate these methods on real-world industrial and robotic datasets, examining the influence of parameters such as the number of masks, model size, and image resolution on the quality of explanations. Our experiments use single-stage object detection models applied to two safety-critical robotic environments: i) a shared human-robot workspace where safety is of paramount importance, and ii) an assembly area of battery kits, where safety is critical due to the potential for damage among high-risk components. Our findings evince that D-Deletion effectively gauges the performance of explanations when multiple elements of the same class appear in a scene, while D-MFPP provides a promising alternative to D-RISE when fewer masks are used.
@@ 1 Introduction
In this context, the field of Explainable AI (XAI) Barredo Arrieta et al. (2020) aims to enhance the interpretability of AI systems by their audience and ultimately, to enhance the user’s trust in the output of AI-based systems. Leaving aside the category of transparent AI models (which are inherently interpretable and do not require any explanations for a user to understand how they work), explainability methods in XAI can be broadly categorized into white-box and black-box approaches. White-box XAI methods require access to the internal workings of the model, such as weights, activations, or gradients (e.g., Grad-CAM Selvaraju et al. (2017)). While these methods can provide powerful insights, they are often limited by their dependence on specific model architectures, making them difficult to generalize across different models and less accessible to users unfamiliar with AI research/tools. In contrast, black-box XAI methods treat the model as an opaque entity, providing explanations based solely on the model’s input-output behavior without requiring any access to its internal components. However, most black-box XAI methods are designed for classification tasks rather than for object detection Ribeiro et al. (2016); Lundberg and Lee (2017); Petsiuk (2018); Ali et al. (2023).
@@
  • Last but not least, we facilitate the broader adoption of the developed techniques for object detection in real-world use cases by releasing the code publicly in a repository: https://github.com/aklein1995/drise_dmfpp_ddeletion.
@@ 2 Related Work
As stated in the introduction, XAI offers insights into the procedure followed by an AI-based system to elicit their outputs, enabling end-users to understand and eventually trust the decisions output by the AI-based system grounded on objective data Ali et al. (2023). To date, the majority of XAI methods are designed for models learned to address classification tasks. For instance, CAM-based methods like GradCAM Selvaraju et al. (2017), GradCAM++Chattopadhay et al. (2018) and Integrated Gradients Sundararajan et al. (2017) quantify and attribute the pixel-wise importance of a given input according to the gradients with respect a target class. Moreover, making use of backpropagation, LRP Montavon et al. (2019) calculates the contribution that a neuron has with neurons in consecutive layers to get relevance scores. In contrast, perturbation-based techniques work by occluding certain parts of the input and analyzing its impact in the predictions. Within this type of techniques, LIME Ribeiro et al. (2016), approximates a NN with an interpretable model; SHAP Lundberg and Lee (2017) assigns importance values to each input feature based on Shapley values; RISE Petsiuk (2018) generates saliency maps by probing the model with randomly masked versions of the input image; and MFPP Yang et al. (2020) generates masks by dividing the input image into multi-scale superpixels. Nonetheless, none of them have been explicitly extended for object detection tasks –with the exception of RISE, which has been adapted for this purpose– although techniques like SHAP can also be utilized for regression problems.
@@ 3 Background
We now proceed by elaborating on key concepts needed to properly understand the details of the proposed D-MFPP technique and the D-Deletion metric that lie at the core of this work. Concretely, we provide fundamentals for object detection models (Section 3.1) and XAI, with a focus on model-agnostic black-box methods to explain the predictions of object detection models (Section 3.2).
@@ 3.1 Object Detectors
Object detectors are crucial components in computer vision tasks, capable of identifying and localizing objects within an image. They can be broadly categorized into single-stage and two-stage detectors.
Single-stage detectors
They directly predict bounding boxes and class probabilities from input images in a single pass. Popular single-stage detectors, such as YOLO Terven et al. (2023), SSD Liu et al. (2016) and RetinaNet Ross and Dollár (2017), treat object detection as a simple regression problem, straight from image pixels to bounding box coordinates and class probabilities. To this end, they produce a dense grid of bounding box proposals and class probabilities in one step. Specifically, YOLO Terven et al. (2023) divides the input image into a grid and predicts bounding boxes and class probabilities for each grid cell. Although this efficiency is beneficial for real-time applications, it often comes at the cost of accuracy when compared to two-stage detectors
@@
These models, among which Faster R-CNN Ren et al. (2017) can be considered to be the most representative one, follow a more complex approach that divides the detection process into two stages. In the first stage, a Region Proposal Network (RPN) generates a set of candidate object proposals (bounding boxes) from the input image. In the second stage, these proposals are refined and classified into different object categories by a second network. This second stage typically involves a more complex network, such as a convolutional neural network (CNN), which performs classification and further refinement of the bounding box coordinates. This two-step process boosts accuracy by allowing for a more refined feature analysis, though it also slows down processing, making two-stage detectors less suited for applications that require high-speed performance.
@@
where $\mathbf{L}_{i}$ defines the bounding box corners $(x^{i}_{1},y^{i}_{1})$ and $(x^{i}_{2},y^{i}_{2})$; $O_{i}\in[0,1]$ refers to the probability that bounding box $L_{i}$ contains an object of any class; and $\mathbf{P}_{i}$ is a vector of probabilities $(p^{i}_{1},\ldots,p^{i}_{C})$ representing the probability that region $\mathbf{L}_{i}$ belongs to each of $C$ classes. Unlike traditional classifiers, which assign a single class label to an entire image, object detectors must handle both classification and localization simultaneously. This dual task, predicting the class and precise location of each object, increases the complexity of making these models interpretable.
@@ 3.2 Explainable Artificial Intelligence (XAI)
Despite the great performance exhibited by object detectors in manifold applications, their adoption in risk-sensitive scenarios is often hindered by a lack of trust and transparency by the user making decisions based on the detections issued by these models. As introduced previously, research on XAI produce techniques and methods that make the behavior and predictions of AI models understandable to humans without sacrificing performance Gunning et al. (2019). To this end, multiple XAI techniques have been proposed, which can be classified into four broad categories Ali et al. (2023):
@@
Among model-agnostic XAI methods, perturbation-based approaches are commonly used due to their simplicity and effectiveness in revealing which parts of the input are most influential for the model’s predictions. Perturbation-based techniques offer a direct way to assess how changes to the input image affect the model’s output. By systematically altering or masking parts of the input image (using masks to generate perturbed samples), these methods allow inferring the importance of different regions based on the model’s input-output behavior.
@@ 4 Materials and Methods
This section describes the industrial robotics use cases in what refers to the datasets (Section 4.1), object detection model (Section 4.2), XAI methods (Section 4.3) and the explanation quality metrics (Section 4.4) considered in our work. The novel XAI technique and quality metrics proposed in this manuscript are also described in Section 4.3.
@@ 4.1 Industrial Robotics Datasets under Consideration
The datasets used in this manuscript have been collected during the course of the ULTIMATE project, https://ultimate-project.eu/, which features two distinct real robotics use cases Kozik et al. (2024). The first dataset, from PIAP https://piap.lukasiewicz.gov.pl/), involves a collaborative workspace where a human and a robotic arm work together. The second dataset, provided by Robotnik https://robotnik.eu/, focuses on a battery assembly area, where a robotic arm assembles components for a battery kit [^2] While the datasets contain a relatively small number of images, this data shortage is typically encountered in real-world industrial scenarios subject to data availability constraints. Nevertheless, in the use cases under considerations the contextual and scene variability is minimal, yielding short-tailed distributions of the objects to be detected. Therefore, the small datasets described in the paper sufficiently capture the relevant features for the specific object detection tasks addressed by the models..
@@ 4.3 Explainability Methods
We evaluate four popular methods for generating visual explanations of black-box models: LIME, RISE, D-RISE, and D-MFPP. The first two methods, LIME and RISE [^4] These XAI methods have been chosen due to their perturbation-based nature, which aligns closely with the methodology followed by the XAI methods D-RISE and D-MFPP proposed in this work. Both D-RISE and D-MFPP generate explanations through perturbations., were originally developed for image classifiers but can be adapted to object detectors, However, they primarily focus on explaining classification aspects and are not capable of addressing localization characteristics. In contrast, D-RISE is one of the first XAI methods specifically designed for object detectors, providing explanations that encompass both classification and localization. Additionally, we extend the existing MFPP method (originally tailored for classifiers) into a version suitable for object detection, which we refer to as D-MFPP. In what follows we briefly describe them, flowing into a description of the proposed D-MFPP approach:
@@ 4.4 Metrics
Evaluating the performance of attribution-based explainability methods for image data involves assessing how well the generated relevance heatmaps highlight important regions of the input image that contribute to the model’s decision. Generally, according to Hedström et al. (2023), explanation quality metrics can be grouped into six categories based on their logical similarity: faithfulness, robustness, localization, complexity, randomization, and axiomatic metrics. In this study, we focus on two of these categories that are particularly relevant to object detection: localization (Section 4.4.1) and faithfulness (Section 4.4.2).
@@ 4.4.2 Faithfulness
  • Minimum Subset: It follows the same logic as Deletion, but instead of determining the AUC, it considers the required number of pixels that make the prediction to change Gevaert et al. (2024). Given the importance score for each pixel ($S_{I,f}$), Min-Subset is defined as the smallest subset of pixels that needs to be removed to change the model’s prediction. Mathematically:
@@ 4.4.3 Proposed D-Deletion and D-Minimal Subset Metrics
The difference between Deletion and D-Deletion is illustrated in Figure 4. This figure highlights how D-Deletion distinguishes between different objects of the same class by incorporating localization information, leading to more refined and accurate explanations ($\downarrow$ AUC in the Figure’s last row) when multiple objects of the same class are detected in an image. For the sake of clarity, we provide the pseudocode of Deletion in Algorithm 1, where the main difference with respect to D-Deletion are lines 10 to 12.
@@ 5 Experiments and Results
  • RQ1: Which XAI method provides the most reliable and insightful explanations for object detection models?
  • RQ2: Does the D-Deletion metric enhance the trustworthiness of XAI outputs when multiple objects of the same class are present in the image?
  • RQ3: How does the mask generation process influence the quality of explanations, particularly when using similarity scores for object detection? How does D-MFPP behave?
  • RQ4: Do different image dimensions impact the explanations generated by XAI methods? Do models of varying sizes (large, medium, small, nano) focus on different regions of the image in their explanations?
@@
RQ1: Comparison between XAI methods
@@
Table 1: Quantitative metrics of LIME, RISE and D-RISE for the Human-Robot dataset. The table presents the performance of each XAI technique in terms of classification (Deletion, D-Deletion, Min-Subset, D-Min-Subset) and localization metrics (PG and EBPG), with scores reported for each class (Human, Gripper) and the overall average. Lower values are better for metrics marked with $\downarrow$, while higher values are better for those marked with $\uparrow$. Bold values indicate the best average scores across all objects, highlighting the best-performing XAI method for each metric. Values highlighted in gray represent the best scores for each object category (Human or Gripper) and should be interpreted vertically.
@@
Table 2: Quantitative metrics of LIME, RISE and D-RISE for the Battery Assembly dataset. The table presents the performance of each XAI technique in terms of classification (Deletion, D-Deletion, Min-Subset, D-Min-Subset) and localization metrics (PG and EBPG), with scores reported for each object (indiv batt, bms a, bms b, unknown object, batt holder) and the overall average. Lower values are better for metrics marked with $\downarrow$, while higher values are better for those marked with $\uparrow$. Bold values indicate the best average scores across all objects, highlighting the best-performing XAI method for each metric. Gray-highlighted values represent the best scores for each object category (indiv batt, bms a, bms b, unknown object or batt holder) and should be interpreted vertically.
@@
RQ2: D-Deletion metric for scenes with multiple objects of the same class
@@
Table 3: Quantitative results for saliency maps generated with various masks techniques (D-Sliding Window, D-RISE, and D-MFPP) in the Human-Robot dataset. Only the average score across all classes is presented. Results are interpreted vertically within each column (each metric represents a separate column). The best mask configuration for each XAI method and for each metric is highlighted in gray, while the overall best-performing XAI method with its optimal mask configuration is shown in bold. .
@@
Table 4: Quantitative results for saliency maps generated with various masks techniques (D-Sliding Window, D-RISE, and D-MFPP) in the Battery Assembly dataset. Results are interpreted vertically within each column (each metric represents a separate column). The best mask configuration for each XAI method and for each metric is highlighted in gray, while the overall best-performing XAI method with its optimal mask configuration is shown in bold.
@@
RQ3: Influence of the Mask Generation Strategy
Influence of the number of masks
RQ4: Image dimension variation and YOLOv8 complexity
Table 5: Explanations comparison for three image dimensions ($720\times 1280$, $360\times 640$, and $180\times 320$) and varying resolutions ($16\times 16$, $8\times 8$, and $4\times 4$) in the Human-Robot dataset.